TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Dan Dubrick
date: 2003-03-31 02:28:00
subject: 3\18 Pt 1 ISS On Orbit Status 18-03-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

19 Mar 2003

FOR OFFICIAL USE ONLY

ISS On-Orbit Status 3/18/03

Part 1 of 2

All ISS systems continue to function nominally, except as noted
previously or below.

CDR Ken Bowersox and FE-2/SO Don Pettit again spent several hours
supporting troubleshooting of the failed ITCS MTL (internal thermal
control system/moderate temperature loop). Resolution of the problem
is temporarily on hold. [During the water filling/priming of the
replacement PPA (pump package assembly) accumulator, the fill
quantity, as shown by the movement of the accumulator bellows, did
not respond to pressure changes in the system as expected. However,
this morning the crew got the green light to press ahead with filling
and subsequent installation of the PPA in the LAB-1 rack. When
powered up afterwards, an unexpected signature was noticed which
appears to indicate a failed-open, i.e., leaking, check valve. The
issue is currently under assessment, and troubleshooting is on hold
until after the Russian software transition is complete. The ITCS
continues in single-loop mode (LTL, low temperature loop), with a
nominally working LTL pump. In case of its failure, leaving the USOS
(U.S. segment) without internal cooling, a contingency plan is in
place for controlled load shedding to reduce heat production, i.e.
power-down of a larger number of RPCMs (remote power controller
modules) in the USOS. In this event, ISS flight attitude would be
moded to "XPOP tumble", until 3/28.]

In the Service Module (SM), FE-1 Nikolai Budarin continued to
reconfigure the Russian OpsLAN network, viz., incorporating the
"telescience" payload into it, supported by tag-up with a ground
specialist. [This involved, first, the installation of a number of
cables to connect the TMO data exchange transformer (unit FL-001) to
onboard power, then linking the TMO to the BRI SmartSwitch, the BSMM
array matching unit laptop and the TVS television subsystem, with
concurrent reconfiguring of the TVS connection of the PK-3 Plasma
Crystal "telescience" payload. Later in the day, Budarin installed a
hard disk drive (HDD) in the TMO FL-001.] 

Transition of the SM's Terminal and Central computers (TVM and TsVM)
to the new software version 7.01 began on schedule. The procedure,
largely controlled by TsUP "launching" pre-loaded cyclograms during
Daily Orbits 14, 15, 1 & 2, started early in the morning and is
planned to take 9-10 hours. [To support the transition, ISS flight
control was moded to CMG-only at 9:10am EST, to be returned to CMG-TA
(thruster assist) at 3:45pm. Systems deactivated for the process
include Vozdukh, Elektron, SKV-1 air conditioner, and BMP
micropurifier. Yesterday, the two computers were restarted
successfully on all three lanes (subsets), running the old vers. 5.04
software, and each is being reloaded today with the new software on
lane #3 first. After a test (verifying checksums), lanes #1 and #2
are then restarted, and an exchange program is launched to execute
the software 7.01 transfer to them, followed by more checkups. Both
TVM and TsVM should be ready for reboot in operational mode (RRP) by
3:20pm EST.] 

USOS GNC (guidance, navigation & control) is working nominally on its
new software patches. GPS (global positioning system) and RGA (rate
gyro assembly) are prime for navigational source data, i.e.,
attitude, attitude rates, and state vector (position, velocity). The
configuration is good to support the 7.01 transition.

In the SM, Nikolai Budarin changed out the CO2 filter unit (BF) in
the IK0501 gas analyzer, an automated system for measuring CO2, O2,
and H2O in the air as well as the flow rate of the gas being analyzed
(replaced last: 2/5).

Later, after obed/lunch (8:00am), Nikolai completed the monthly
preventive maintenance on the SM air ventilation system, today
cleaning Group A fans and grilles.

Don Pettit continued his support of the ongoing MSG (Microgravity
Science Glovebox) troubleshooting. All E-Box connectors have been
remated, but identification of the fault remains elusive. The next
step, starting today, is restoration of MSG to its nominal
operational configuration, pausing between steps for ground
commanding/evaluation. This will continue for the next few days as
time permits around the ITCS/MTL troubleshooting. [Today's activities
included a laptop reconfiguration, followed by powered facility runs
with sequential adding of MSG MLC (MSG laptop computer) and a powered
video drawer. Seven short-duration on-board activities were planned,
with associated ground commanding between each on-board activity.]

The next Foot (foot/ground reaction forces during space flight)
experiment activities have slipped by a day due to the MTL/PPA R&R.
EMG (electromyography, i.e., muscular electric signals) calibration
for Foot and the subsequent day-long experiment run is scheduled for
tomorrow, and Ken Bowersox today unstowed, set up and prepared the
payload equipment for it. 

Pettit conducted the daily routine maintenance of the SM's SOZh life
support system, including ASU toilet facilities, while Budarin
prepared the daily IMS inventory delta file.

Nikolai performed his daily status check on the Russian BIO-5
Rasteniya-2/Lada-2 ("Plants-2") plant growth experiment in the SM.

All crewmembers completed their daily physical exercise program on
RED (resistive exercise device), TVIS (treadmill with vibration
isolation and stabilization), and VELO with load trainer. Don
performed the weekly routine maintenance on the treadmill, and Sox
conducted the bi-weekly inspection of the RED after his resistive
exercise session. CEVIS (cycle ergometer with vibration isolation)
control panel continues failed. 

At 10:40am, Bowersox and Pettit fielded media questions in two live
TV press interviews of about 10 minutes each, with WISH-TV,
Indianapolis, IN, and the Silverton, OR, Appeal Tribune.

Lab ppCO2 (carbon dioxide partial pressure) level this morning was
4.2 mmHg. Vozdukh CO2 scrubber in the SM is deactivated for the
duration of the Russian software transition. CDRA (CO2 removal
assembly) in the Lab was later powered up by MCC-H.

Vozdukh's failure last Friday (3/14) was due to a damaged power
cable, as suspected. After repair by the crew (there was no spare on
board), the system was back up in operation on Saturday. A set of
spare cabling is now being prepared for launch on 11P/Progress.

Moscow has reported the results of the two Progress thruster dynamic
tests last week. The first test reboost on 3/12, with the #2 thruster
manifold, was perfectly nominal. However, the firing on 3/14 of the
suspected #1 manifold showed the same abnormal signature as it did
during the original reboost on 2/11 (ref. also Status Report 2/14).
[The burn spent nearly equal amounts of fuel and oxidizer (8 kg ea.),
indicating an oxidizer underuse of 30% from the nominal. Thrust was
only 70% of nominal. There will be a report by the investigating
Commission. As of now, the major reboost in April will be on
manifold #2 only, with #1 to be used solely in an emergency. The burn
will use up all remaining Progress fuel (unsymmetrical
dimethylhydrazine, UDMH), but it is still to be decided whether the
left-over oxidizer (nitrogen tetroxide, N2O4) will be transferred to
the ISS or jettisoned with the vehicle.]

The Stage EVA2 by Bowersox and Pettit is now set for 4/8 (Tuesday).
On the day before, the GNC MDM will receive a software load to
support one of the tasks planned for the spacewalk (the CMG patch
panels reconfiguration at the Z1 truss).

 - Continued -

@Message posted automagically by IMTHINGS POST 1.30
--- 
* Origin: SpaceBase(tm) Pt 1 -14.4- Van BC Canada 604-473-9358 (1:153/719.1)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.